# Geo4D: Leveraging Video Generators for Geometric 4D Scene Reconstruction

### v1

@@ 5 Discussion and Conclusion
We have introduced Geo4D, a novel approach that adapts a video generator to dynamic 4D reconstruction. By building on a pre-trained video generator, Geo4D achieves excellent generalization to real data despite being trained only on 4D synthetic data. We have also shown the benefits of predicting multiple modalities and fusing them at test time via optimization. Our model outperforms state-of-the-art methods on video depth and camera rotation prediction, particularly on challenging dynamic scenes.
Despite these successes, our approach has limitations, one of which is that the point map encoder-decoder is still not entirely accurate, and that, in turn, is a bottleneck on the overall reconstruction quality.
Our approach also opens a path to integrating 4D geometry in video foundation models, e.g., to generate 3D animations from text, or to provide a more actionable signal when the video model is used as a proxy for a world model.
@@ Acknowledgments.
The authors of this work were supported by Clarendon Scholarship, ERC s-UNION, and EPSRC EP/Z001811/1 SYN3D.
@@
In this supplementary document, we provide additional materials to supplement our main submission. In the supplementary video, we show more visual results using our method. The code will be made publicly available for research purposes.
@@ 6.1 Training dataset
As shown in Tab. 5, we use five synthetic datasets for training: Spring [50], BEDLAM [2], PointOdyssey [119], TarTanAir [93], and VirtualKitti [6]. Although all datasets we used are synthetic data sets, we found that there are still some missing depth pixels in PointOdyssey [119]. So we apply max pooling to inpaint those missing pixels. During training, we sample each data set with the corresponding ratio in Tab. 5. For each sample, we sample 16 frames from the sampled sequence with the sampling stride randomly chosen from $\{1,2,3\}$ to allow our diffusion model to adapt to different input videos with various FPS.
@@ 6.2 Optimization details
Algorithm 1 Multi-Modal Alignment Optimization
1:$\bm{X}^{i,g},\bm{D}^{i,g},\bm{r}^{i,g}\leftarrow$ Predicted by our diffusion model
2:$\bm{D}^{i}_{\text{p}},\lambda^{g}_{\text{p}},\bm{R}^{g}_{\text{p}},\bm{\beta}^{g}_{\text{p}}\leftarrow$ Initialized by Umeyama algorithm
@@
8:          $\lambda^{g}_{\text{d}},\beta^{g}_{\text{d}}\leftarrow\arg\min\mathcal{L}_{\text{d}}$q. 5)
9:          $\bm{R}^{g}_{\text{c}},\lambda^{g}_{\text{c}},\bm{\beta}^{g}_{\text{c}}\leftarrow\arg\min\mathcal{L}_{\text{c}}$ec. 3.3)
10:    else if Iteration $<$ Align start iteration then
@@
6.3 Ablation study for number of denoising steps
@@
Abs Rel $\downarrow$ $\delta<1.25$uparrow ATE $\downarrow$ RPE trans $\downarrow$ RPE rot $\downarrow$
@@
We study the influence of the number of denoising steps during inference. As shown in Tab. 6, the model achieves optimal performance around steps equal to 5. Compared to the video generation task, where a larger denoising step usually produces a more detailed generated video, 4D reconstruction is a more deterministic task, which requires fewer steps. Similar phenomena are also observed in [22], which utilize a video generator for the video depth estimation task.
Figure 5: Additional qualitatively results. Our method generalizes well to various scenes with different 4D objects and performs robustly against different camera and object motion.
@@ 6.4 Ablation study for fine-tuned point map VAE
Abs Rel $\downarrow$ $\delta<1.25$uparrow ATE $\downarrow$ RPE trans $\downarrow$ RPE rot $\downarrow$
@@
As stated in the main paper, we added an additional branch to predict the uncertainty for our point map VAE and fine-tuned it based on Eq. 3. We perform the ablation study on our fine-tuning strategy. As shown in Tab. 7, our fine-tuned point map VAE achieves consistent better performance on both video depth estimation and camera pose estimation tasks compared with the original pre-trained image VAE, demonstrating the necessity and the effectiveness of our fine-tuning strategy.
@@ 7 Visualization
As shown in Fig. 5, we show additional visualization results for both the indoor, outdoor, and driving scene. Although our model is only trained on synthetic datasets, it is generalized to real-world data with diverse objects and motions.

### v2

@@ 5 Discussion and Conclusion
We have introduced Geo4D, a novel approach that adapts a video generator for dynamic 4D reconstruction. By building on a pre-trained video generator, Geo4D achieves excellent generalization to real data despite being trained only on synthetic 4D data. We have also demonstrated the benefits of predicting multiple modalities and fusing them at test time via optimization. Our model outperforms state-of-the-art methods on video depth and camera rotation prediction, particularly in challenging dynamic scenes.
Despite these successes, our approach has limitations. One is that the point map encoder-decoder is still not entirely accurate, which in turn is a bottleneck for the overall reconstruction quality.
Our approach also opens a path to integrating 4D geometry into video foundation models, e.g., to generate 3D animations from text, or to provide a more actionable signal when the video model is used as a proxy for a world model.
@@ Acknowledgments.
The authors of this work were supported by Clarendon Scholarship, ERC 101001212-UNION, and EPSRC EP/Z001811/1 SYN3D.
@@
In this supplementary material, we provide additional information to supplement our main submission. The code is available here for research purposes: github.com/jzr99/Geo4D
@@ 6.1 Training Dataset
As shown in Tab. 5, we use five synthetic datasets for training: Spring [50], BEDLAM [2], PointOdyssey [119], TarTanAir [93], and VirtualKitti [6]. Although all datasets are synthetic, we found that some depth pixels are missing in PointOdyssey [119]. To address this, we apply max pooling to inpaint the missing pixels. During training, we sample each dataset according to the ratios in Tab. 5. For each sample, we select 16 frames from the sequence, with the sampling stride randomly chosen from $\{1,2,3\}$ to allow our diffusion model to adapt to input videos with various frame rates.
@@ 6.2 Optimization Details
The overall optimization process is outlined in Algorithm 1. We first predict all three modality maps using our diffusion model for each video clip $g$. The predicted point maps are then roughly aligned based on the overlapping frames using the Umeyama algorithm [84]. The camera intrinsic $\textbf{K}^{k}$ is initialized by minimizing the projection error of the point map $\bm{X}^{k,g^{k}}$ in its reference (first) frame $k$ within each window group $g^{k}$. The camera extrinsics are then initialized using the RANSAC PnP algorithm. In the first stage of optimization, the point maps are roughly disentangled into camera pose and depth map. The disparity map is then aligned with the global depth inferred from point maps by solving Eq. 5 from the main paper to obtain the scale and shift parameters. The camera parameters extracted from the predicted ray map are aligned with the global camera trajectory based on the reference (first) frame of each video clip $g$ via Sec. 3.3 from the main paper. After initializing all the alignment learnable parameters, including rotation $\bm{\mathbf{R}}^{g}_{*}$, scale $\lambda^{g}_{*}$, and shift $\bm{\mathbf{\beta}}^{g}_{*}$ across different modalities, where $\in\{\text{p},\text{d},\text{c}\}$, we jointly optimize all the learnable parameters by Eq. 10. Specifically, we set the weights for each loss term in Eq. 10 as $\alpha_{1}=1,\alpha_{2}=2,\alpha_{3}=0.005,\alpha_{4}=0.015$ to roughly equalize the scale of the different losses.
Algorithm 1 Multi-Modal Alignment Optimization
@@
[…]
@@
8:     $\lambda^{g}_{\text{d}},\beta^{g}_{\text{d}}\leftarrow\arg\min\mathcal{L}_{\text{d}}$Eq. 5)
9:     $\bm{R}^{g}_{\text{c}},\lambda^{g}_{\text{c}},\bm{\beta}^{g}_{\text{c}}\leftarrow\arg\min\mathcal{L}_{\text{c}}$Sec. 3.3)
10:  else if Iteration << Align start iteration then
@@
7 Additional Analysis
7.1 Ablating the Number of Denoising Steps
@@
Abs Rel $\downarrow$ $\delta<1.25$ $\uparrow$ ATE $\downarrow$ RPE trans $\downarrow$ RPE rot $\downarrow$
@@
We study the influence of the number of denoising steps during inference. As shown in Tab. 6, the model achieves optimal performance after around 5 steps. Compared to the video generation task, where a larger number of denoising steps usually produces a more detailed generated video, 4D reconstruction is a more deterministic task, which requires fewer steps. Similar phenomena are also observed in [22], which uses a video generator for video depth estimation.
Figure 5: Additional qualitative results. Our method generalizes well to various scenes with different 4D objects and performs robustly against different camera and object motions.
@@ 7.2 Ablation Study for Fine-Tuned Point Map VAE
Abs Rel $\downarrow$ $\delta<1.25$ $\uparrow$ ATE $\downarrow$ RPE trans $\downarrow$ RPE rot $\downarrow$
@@
As stated in the main paper, we added an additional branch to predict the uncertainty for our point map VAE and fine-tuned it based on Eq. 3. We perform an ablation study on our fine-tuning strategy. As shown in Tab. 7, our fine-tuned point map VAE achieves consistently better performance on both video depth estimation and camera pose estimation tasks compared with the original pre-trained image VAE, demonstrating the necessity and effectiveness of our fine-tuning strategy.
7.3 Analysis of Multi-Modal Representation
Point maps (PMs) and disparity maps (DMs) are complementary. DMs better represent near objects, while PMs are more depth-agnostic (e.g., human vs house in Fig. 6 (b,c)). As in prior work, DMs are affine invariant (which here makes them range-compatible with the pretrained RGB VAE); their scale and shift, needed to recover undistorted geometry, are inferred by matching them to the predicted PMs. Ray maps (RMs) help infer the camera pose when PMs fail to represent points at infinity (such as the sky in Fig. 6 (e)). We observed that PMs tend to be noisier than DMs, so we prioritized modeling the PMs’ uncertainty. Per-pixel uncertainty for ray maps are less meaningful given the high degree of correlation between individual rays. During multi-modal alignment, we align global point clouds with DMs in disparity space and with PMs in linear space. This naturally gives more weight to near points, which tend to be estimated well by DMs, and weighs points based on uncertainty with PMs, thus taking advantage of both modalities.
Figure 6: Visualization of different geometric modality maps.
@@ 8 Visualization
Figure 5 shows additional visualizations for indoor, outdoor, and driving scenes. Although our model is only trained on synthetic datasets, it generalizes to real-world data with diverse objects and motions.
9 Limitations
Although our method performs well and generalizes to a wide range of in-the-wild videos, it can struggle in cases involving significant changes in focal length or extreme camera motion throughout a sequence. This limitation likely stems from the lack of focal length variation in our training data. Incorporating more sequences with diverse camera movements and zooming effects could help mitigate this issue. Moreover, due to the inherent temporal attention mechanism in our network architecture, our approach currently supports only monocular video input. Extending the method to handle multi-view images or videos is a promising direction for future work.